I am trying to get ahead on my homework. How do you find the focus point of y^(2)=(8/3)x? Because from there I have to find the latus rectum. I know how to do that part from the focus point. Thank you!!
Answer by lwsshak3(11628)

I am trying to get ahead on my homework. How do you find the focus point of y^(2)=(8/3)x? Because from there I have to find the latus rectum. I know how to do that part from the focus point.
***
Give data shows parabola opens right.
Its basic equation: y^2=4px
For given parabola: y^(2)=(8/3)x
4p=8/3
p=2/3
vertex: (0,0)
axis of symmetry: y=0 or x-axis
focal point=(2/3,0) (p-distance to right of vertex on axis of symmetry.
latus rectum=4p=8/3